• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 07:46:21 on July10,2000

    Entry number 45478

    keyword=Owl Shift Summary

    00:00 Still no beam in Hall A. The spin dance is supposed to begin after the
    beam comes to the Hall. BPM problems are not fixed yet.
    00:30 Beam in Hall A (0.7uA). 12C target. Seems the BPMs are working now.
    The spot ++ shows rastered beam spot picture. Joe came to prepare for the compton
    01:30 Eugene came to do the moller measurement.
    04:00 Start moller measurement